OpenStudy (anonymous):

which ratio is proportional to 3/5 a) 9/10 b)12/20 c)15/20 d)5/7

OpenStudy (anonymous):

there are infinite ratios to which are proportional to 3/5

OpenStudy (anonymous):

3/5 =3*4/5*4= 12 / 20